OpenStudy (anonymous):

3)Let \(f\) differentiable on the open interval \(I\) and let \(f^{'}\)be bounded on \(I\). a) Show that \(|f(x)-f(y)| \leq |x-y| sup |f^{'}(t)| \) holds for all \(x, y \in I, x < y\)

OpenStudy (anonymous):

By the mean value theorem there is c in I such tthat \[f(x) - f(y)= f'(c)(x-y)\\ |f(x) - f(y)|=| f'(c)||(x-y)| \le M |x-y|\\ \] Where M is the bound of f' on I.

OpenStudy (anonymous):

thank you very much Mr Elias

OpenStudy (anonymous):

yw